Circular No. 8159

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gh IN NGC 2466 Further to IAUC 8137, L. A. G. Monard reports his discovery of an

apparent supernova on unfiltered CCD images taken on June 29.737 (mag ∼ 15.7 ± 0.3) and 30.694 UT (mag 15.6). SN 2003gh is located at α = 7^{h}45^{m}17^{s}.93, δ = −71^{o}24′37′′.3 (equinox 2000.0), which is 6′′ west and 0′′.5 south of the nucleus of NGC 2466. The new object is not present on the Digitized Sky Survey (limiting red mag 19) or on an image taken by Monard on June 7.707 (limiting red mag 18.5).

SUPERNOVAE 2001co, 2003H, 2003dg, AND 2003dr A. V. Filippenko, R. Chornock, and B. Swift, University of Califor-

nia, Berkeley; M. Modjaz, Harvard-Smithsonian Center for Astrophysics; R. Simcoe, California Institute of Technology; and M. Rauch, Carnegie Observatories, report that inspection of CCD spectra (range 320–1000 nm) obtained on 2001 June 29, July 16, and 27 UT with the Shane 3-m telescope at Lick Observatory, and on 2001 Aug. 17 with the Keck-I telescope (+ LRIS), reveal that SN 2001co (IAUC 7643) is a peculiar type-Ib or type-Ic supernova. Lines of H and He are absent, but the Ca II near-infrared triplet is very strong in emission, and the [Ca II] doublet near 730 nm is also present. No other clear emission or absorption features are visible in the combined, noisy spectrum. The spectral characteristics of this ‘calciumrich supernova’ were unprecedented, and perhaps suspect (given the low signal-to-noise ratio), until Keck-I observations of SN 2003H (IAUC 8045, 8046), obtained on 2003 Feb. 28 UT by Filippenko and Chornock, revealed a very similar-looking spectrum: the only strong features are the [Ca II] and especially the near-infrared Ca II emission lines. SN 2003H had been classified as type Ib by Hamuy (IAUC 8047, 8049), but the Keck spectrum ∼ 2 months after explosion is clearly very unusual; in normal type-Ib and type-Ic supernovae, prominent [O I] emission emerges in the nebular phase. Filippenko and Chornock add that inspection of CCD spectra (range 320– 990 nm) obtained on June 29 with the Keck-I telescope (+ LRIS) reveals that SN 2003dg (IAUC 8113) and SN 2003dr (IAUC 8117) are peculiar type-Ib or type-Ic supernovae of the ‘calcium-rich’ variety described above, but evidently at a later phase than seen for SN 2001co and SN 2003H because the [Ca II] emission is stronger than the near-infrared Ca II triplet. [O I] 630.0- and 636.4-nm emission is much weaker than in normal type-Ib and type-Ic supernovae many months after the explosion.

2003 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green